OpenStudy (dtan5457):

Help me graph -cos(x+pi/2)-2

OpenStudy (dtan5457):

@mathstudent55

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

does it help to know that \[\cos(x+\frac{\pi}{2})=-\sin(x)\]?

OpenStudy (misty1212):

should right? all you have to do is graph \[y=\sin(x)\] shifted down two units

OpenStudy (dtan5457):

in other words the reflection of sin is cosine, vice versa?

OpenStudy (misty1212):

no in other words ... \[\cos(x+\frac{\pi}{2})=-\sin(x)\] that is all

OpenStudy (misty1212):

if you shift cosine \(\frac{\pi}{2}\) units to the left, you get \(-\sin(x)\)